a airplane travels 200 km straight course at an angle 22 degrees due east of north it then changes course by moving 80 km north before reaching its destination. calculate the resultant displacement of the airplane using component method.
1
Expert's answer
2021-09-14T14:15:10-0400

"s_x=200\cdot\sin22°=74.9\ (km)"


"s_y=80+200\cdot \cos22°=265.4\ (km)"


"s=\sqrt{74.9^2+265.4^2}=275.8\ (km)"